OpenStudy (anonymous):

how would you solve 6x-3y=15 and 7x+4y with elimination?

OpenStudy (tkhunny):

I would use elimination! :-) 6x-3y=15 7x+4y ----- There seems to be something missing.

OpenStudy (anonymous):

oh whoops. 7x+4y=1o is the second equation

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

6x-3y = 15 7x+4y = 10 Multiply the first by 4 24x - 12y = 60 Multiply the second by 3 21x + 12y = 30 Anything strike you, yet?

OpenStudy (anonymous):

thank you so much! that was the only problem i didnt quite understand on my homework

OpenStudy (anonymous):

x=2 y=-1